1.Interest rate is critical to the speed at which your investment grows. If $1 is invested at 2%, it takes approximately 34.9 years to double. If $1 is invested at 5% it takes approximately 14.2 years to double. How many years will it take for $1 to double if invsted at 10%; and 12%? It asks to use the future value table or compound amount of $1.00.
-My answer for 10% would be 7.1 years since it is half of 14.2 but more than 5%? I have no idea what formula to use. Have no idea how I would even get these answered? What formula should be used if I have no idea of time?
2. At what interest rate would you need to invest to have your money double in 10 years? same it asks to use the table and use the 10 year row that represents doubling? I'm lost!

Answer by rfer(16322)

here is when the rule of 72 works nicely, but it isn't exact.
divide 72 by the interest rate as a whole number not a decimal.
72/10%=7.2 yrs
72/12%=6 yrs
----------------
72/10yrs=7.2%